OpenStudy (anonymous):

For an angle Θ with the point (12, −5) on its terminating side, what is the value of cosine?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 ???

OpenStudy (anonymous):

Anyone

OpenStudy (anonymous):

Anyone at all

OpenStudy (anonymous):

find the hypotenuse via pythagoras

OpenStudy (anonymous):

I am telling you right now it is not -5/13

OpenStudy (anonymous):

then it will be \[\cos(\theta)=\frac{12}{13}\]

OpenStudy (anonymous):

yeah, that is the sine!

OpenStudy (anonymous):

cosine is the first coordinate over the hypotenuse since is the second

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Ok, thank you

OpenStudy (anonymous):

yw